OpenStudy (anonymous):

solve the equation 10-5b=-8-14b -2/19 -2 2/19 18/19

OpenStudy (anonymous):

what equation??

OpenStudy (anonymous):

The solution is the second one listed above, b = -2

hero (hero):

Hint: Put like terms on the same side: 10 + 8 = -14b + 5b =

hero (hero):

@robtobey , you should read the code of conduct. Members of OS are not allowed to give answers anymore.

OpenStudy (anonymous):

thank you robtobey

hero (hero):

Did he really help you? Do you know how to solve the problem now? Or did he just give you the answer outright?

hero (hero):

Anti-Cheating Policy What is OpenStudy's Anti-Cheating Policy? Openstudy.com is proud to be a pioneer in the online education community and is dedicated to advancing knowledge, understanding and participation in student success. Through OpenStudy.com, students, educators and subject enthusiasts from around the globe participate in a dynamic social learning experience that stands above the rest. The resources we provide, study rooms and users, are here to help you study smarter and study with friends. We believe you will actually enjoy learning now! We work diligently to prevent misuse, but ultimately, you make the choice to use or misuse the resources we provide. We require every member to abide by OpenStudy’s Anti-cheating Policy. Openstudy.com will take swift action against anyone found misusing the community’s site. In addition to refunding any membership fee, we will ban users who cheat from the site permanently. A message to students: Please understand that as a member of the OpenStudy community you agree to abide by the academic policies or honor codes of your school and course, and report violations promptly to us. Please keep in mind that Openstudy is a community for educators as well as students.

OpenStudy (anonymous):

yes, because i needed to know the answer to know how to do the first step of the problem. So yes he helped me

hero (hero):

Giving the answer does not help you unless you can demonstrate that you know how to solve it now. When you actually can actually post the steps, then I'll believe you. Either way, users are not supposed to post just answers only.

hero (hero):

@Tigeris97, you can post your steps at any time.

OpenStudy (anonymous):

fine! i will

hero (hero):

Remember, the directions say to solve the equation.

OpenStudy (anonymous):

10-5b=-8-14b +14 +14 10+9b=-8 -10 -10 9b=-18 b=-2 Boooooooooom!

hero (hero):

How do I know robtobey didn't private message you? Plus it took you too long.

OpenStudy (anonymous):

@Hero As a minor contributor for over a year now, beginning when Openstudy operated under another site name, with only 1200 students participating, no medal awards, and no "regulations" regarding this and that, the atmosphere was entirely different and rewarding. Now with the addition of what appears to be a cadre of well meaning solution police/management, it is probably time for me to bail out.

OpenStudy (anonymous):

It took long because i had to figure out the steps and im 1 day new to openstudy so i didnt know there was a private message thing

hero (hero):

That's your choice @robtobey , I'm just alerting you to the policy change. With as long as you've been on here, you should want to help students, not just give them the answer. Giving answers is analogous to cheating.

OpenStudy (anonymous):

wellllllllllllll

OpenStudy (anonymous):

i got the answer ad u feel salty so case closed! caso cerado !

hero (hero):

I suppose you didn't read the anti-cheating policy I posted above. It clearly states that you can be permanently banned for posting questions with the intent of receiving answers only.

hero (hero):

Many people don't know apparently. That's the point of the posts I made. To let people know that such policies are in place.